a bag contains 4 white ball,5 red ball,6 green balls 3 balls are drawns at random.wht is the probabilty that a white ,ared ,gerrn ball is drawn
Found 3 solutions by richwmiller, drk, stanbon:
Answer by richwmiller(17219)

Your spelling is horrible.
You don't state how the balls are drawns, as you say.
All at once or one after another with or without replacement.

Answer by drk(1908)

P(G, R, and W)
Probability is success / total
successful ways are 6c1, 5c1, and 4c1
the total is (15c3) = 455
So,
P(G, R, and W) = 6*5*4 / 455 = .2637

Answer by stanbon(75887)

A bag contains 4 white ball,5 red ball,6 green balls 3 balls are drawns at random.what is the probabilty that a white ,a red ,a green ball are drawn.
-----
# of possible outcomes: 15C3 = 455 sets of 3 balls
----------------------------
# of all white = 4C3= 4
# of all red = 5C3 = 10
# of all green = 6C2 = 15
-----------------------------
# of 2W and 1red = (4C2)*5 = 30
# of 2W and 1green= (4C2)*6 = 36
# of 2R and 1W = (5C2)*4 = 40
# of 2R and 1G = (5C2)*6 = 60
# of 2G and 1red = (6C2)*5 = 75
# of 2G and 1W = (6C2)4 = 60
---------------------------------
# of ways to get 1R,1W,1G = 455-[29+226+75] = 125
===========
P(1R,1W,1G) = 125/455 = 0.2747
====================================
Cheers,
Stan H.